OpenStudy (anonymous):

Exact Value: cos(π/12)

OpenStudy (anonymous):

I've started heading in the direction of cos(180/12) but I am not sure what is the best way to break this down.

OpenStudy (anonymous):

cos(180/12) = cos(15)

OpenStudy (shamim):

ya u can do in this way

OpenStudy (shamim):

\[\cos15=\cos(45-30)=\cos45\cos30+\sin45\sin30\]

OpenStudy (anonymous):

Okay I see what you did

OpenStudy (anonymous):

Difference Identity Formula for cos

OpenStudy (shamim):

\[=(1/\sqrt{2})(\sqrt{3}/2)+(1/\sqrt{2})(1/2)\]

OpenStudy (shamim):

\[(\sqrt{3}/2\sqrt{2})+(1/2\sqrt{2})\]

OpenStudy (shamim):

\[(\sqrt{3}+1)/2\sqrt{2}\]

OpenStudy (calculusfunctions):

@shamim we don't just give out answers. We teach and guide the student to the correct answer.

OpenStudy (shamim):

ok

OpenStudy (shamim):

did u get my solution

OpenStudy (anonymous):

@calculusfunctions I wasn't even paying attention after I knew which formula to use lol

OpenStudy (anonymous):

I got:

OpenStudy (shamim):

thanks a lot

OpenStudy (anonymous):

\[\frac{ √(2+√3) }{ 2 }\]

OpenStudy (anonymous):

(No offense intended @shamim )